OpenStudy (anonymous):

Write this ratio $17.20 to $5.04 as a fraction in lowest terms

OpenStudy (solomonzelman):

\[\frac{17.20}{5.04}\]\[\frac{17.20\color{red} {\times 100 }}{5.04\color{red} {\times 100 }}\]\[\frac{1720}{504}\]\[\frac{1720\color{red} {\div 8}}{504\color{red} {\div 8}}~~~->~~~~\frac{215}{63}\]

OpenStudy (pixiedust1):

nice one solomon!

OpenStudy (anonymous):

Thank you Solomon.